OpenStudy (anonymous):

exact value using sum or difference identity ; sin255 degrees

pooja195 (pooja195):

Look for a combination of "exact value" angles that gives you 255... how about 120 + 135 so now sin (255) = sin (120 + 135)= sin 120 cos 135 + cos 120 sin 135 = (sqrt3/2)(-sqrt2/2) + (-1/2)(sqrt2/2) = -sqrt6/4 - sqrt2/4 = -(sqrt6 + sqrt2)/4 or, rearranging to match your choices: (-sqrt2 - sqrt6)/4 (a)

OpenStudy (campbell_st):

this is really the sum of 120 degrees and 135 so you would have sin(120 + 135) = sin(120)cos(135) - cos(120)sin(135) both angles have exact values....

OpenStudy (anonymous):

Thank you so much!!!!